OpenStudy (anonymous):

can someone check if my answer's right? fan and medal! Without solving, how many solutions does the following equation have? Justify your answer in your written work. \begin{cases} -4y = -8x^{2}+80x-204 \\ y = 2x^{2}+12x+20 \end{cases}

OpenStudy (anonymous):

-4y = -8x^2 + 90x -204 y = 2x^2 -20x -51 2x^2-20x-51 = 2x^2 +12x +20 0=32x+71

OpenStudy (anonymous):

so how many solutions will this have? :P

OpenStudy (anonymous):

2x^2 -20x + 51, not - 0= 32x - 31 one solution.

OpenStudy (anonymous):

oh psht haha, thank you so much! :DD

OpenStudy (anonymous):

Anytime :)